Quasar, Inc. sells clothing, accessories, and personal care products for men and women through its retail stores. Quasar reported the following data for two recent years:

   Year 2    Year 1
Sales $3,228,060 $3,321,500
Accounts receivable 244,550 237,250

Assume that accounts receivable were $273,750 at the beginning of Year 1.

a. Compute the accounts receivable turnover for Year 2 and Year 1. Round to one decimal place.

Year 2:
Year 1:

b. Compute the days' sales in receivables for Year 2 and Year 1. Round interim calculations and final answers to one decimal place. Use 365 days per year in your calculations.

Year 2: days
Year 1: days
Answer #1

Accoounts Receivable Turnover Ratio = Net Credit Sales/Average Accounts Receivables

Average Accounts receivables= (Openind Receivable + closing receivables) / 2

Year 1

Average Receivables = ($273,750 + $237,250) / 2 = $255,500

Net Credit Sales = $3,321,500

Accounts Receivable Turnover Ratio = $3,321,500 / $255,500 = 13 Times

Note: Sales here are assumed to be on credit wholly.

Note: Closing receivables of one year becomes opening of another year.

Year 2

Average Receivables = ($237,250 + $244,550) / 2 = $240,900

Net Credit sales = $3,228,060

Accounts Receivable Turnover Ratio = $3,228,060 / $240,900 = 13.4 Times

Days Sale In Receivables = No. of Days in a Year / Accounts Receivable Turnover Ratio

Year 1

Days Sale In Receivables = 365 / 13 = 28.07 i.e. 28 days

Year 2

Days Sale In Receivables = 365 / 13.4 = 27.23 i.e. 27 days